# IEEEabrv.bib

### v1

@@ Abstract
We introduce SPOT, an object-centric imitation learning framework. The key idea is to capture each task by an object-centric representation, specifically the SE(3) object pose trajectory relative to the target. This approach decouples embodiment actions from sensory inputs, facilitating learning from various demonstration types, including both action-based and action-less human hand demonstrations, as well as cross-embodiment generalization. Additionally, object pose trajectories inherently capture planning constraints from demonstrations without the need for manually crafted rules. To guide the robot to execute the task, the object trajectory is used to condition a diffusion policy. We show improvement compared to prior work on RLBench simulated tasks. In real-world evaluation, using only eight demonstrations shot on an iPhone, our approach completed all tasks while fully complying with task constraints. Project page: https://nvlabs.github.io/object_centric_diffusion
@@ I Introduction
Learning from demonstrations is an effective method for acquiring complex robotic manipulation skills. This approach simplifies learning vision-based manipulation policies by training directly on demonstration data, thus avoiding the need for intricate hand-crafted rules or laborious robot self-exploration in real-world settings. However, creating a structural representation that captures task-relevant information, remains resilient to environmental perturbations, and generalizes across the robot embodiment remains a significant challenge.
Previous work [reed2022generalist, jang2022bc, ahn2022can] has utilized raw RGB inputs for end-to-end visuomotor policy training. However, the absence of explicit 3D representation makes it difficult to generalize to various viewing angles and out-of-distribution spatial configurations. To address these issues, RVT [goyal2023rvt] and PerAct [shridhar2023perceiver] use the entire 3D scene as input, represented through point clouds and 3D volumes. While these methods consider the overall 3D scene structure, they include excessive redundant information, reducing data efficiency and necessitating extensive demonstration data collection. Additionally, the tight coupling of observation and action pairs complicates deployment across robotic embodiments, camera setups, or learning from action-less data like human hand demonstrations.
Object-centric methods address these problems by extracting object-centric information from visual observation, such as object detection [zhu2023viola], 6D object poses [wen2022you], 2D or 3D flows [zhu2024vision, xu2024flow, yuan2024general], to serve as inputs for downstream policies. By leveraging visual foundation models, these methods exhibit impressive generalization capabilities in new environments and demonstrate strong data efficiency, often requiring only a minimal amount of demonstration. Nevertheless, most works do not model the holistic object pose trajectories or focus only on the last-inch manipulation policy. This oversight necessitates extra human effort to manually create rules for constrained planning, especially in tasks where intermediate actions are crucial for success. For example, in tasks like pouring water or serving a plate, the container must remain upright throughout to prevent spillage. Such constraints are non-trivial to automatically learn with existing object-centric frameworks.
@@
To address these issues, we introduce SPOT, an object-centric diffusion policy framework that combines diffusion policy and object-centric representation for manipulation. As shown in Fig. 1, it uses the SE(3) object-centric trajectory relative to the target as the representation. This approach allows us to decouple the embodiment action from sensory input, making it easier to learn from various types of demonstration data, whether they involve action-based or action-less human hand demonstrations. In particular, opposite to the standard diffusion policy [chi2023diffusion, ze20243d] which directly produces robot end-effector action, our diffusion model learns to synthesize target SE(3) object-centric pose trajectories along the horizon, conditioned on the current observed object pose. The training data is obtained by extracting such object-centric information from demonstration videos. For robot execution, the action is computed from the task space to transport the object following the target synthesized trajectory. By recurrently switching between future trajectory synthesis and trajectory following, we achieve closed-loop control with enhanced robustness under dynamic uncertainty, such as in-hand object motion during grasping or placement. Unlike previous methods that focus solely on the last inch of manipulation, our approach trains on the pose trajectory data of the complete horizon, eliminating the need for manually crafted global planning rules before the last-inch stage. It also leverages the data scaling law when large amounts of demonstration data are available while maintaining data efficiency in few-shot demonstration scenarios.
We evaluate the approach on RLBench [james2020rlbench], where it significantly outperforms in high-precision and long-horizon tasks in a challenging single camera setting. In real-world tasks, with just 8 demonstrations shot on an iPhone, the resulting policies complete all tasks while fully complying with task constraints. We will make all datasets and code available so others can reproduce and expand upon this work. In summary, our framework has the following desired properties:
@@
    Flexibility on diverse demonstration data. Our framework can learn from action-based robotic demonstrations and action-less data from human demonstrations. This flexibility enables easier data collection using accessible setups, such as a hand-held iPhone, without the need for robotic arms, as our experiments demonstrate.
@@
    Autonomy on complex task constraints. Our framework considers the entire sequential object pose trajectory. This approach allows it to automatically learn planning constraints from demonstrations, eliminating the need for manually crafted rules.
@@
    Generalization. Our object-centric representation allows us to ignore the task-irrelevant information from the background and thus generalizes robustly to varying lighting conditions, scene setup, and sensory configurations.
@@
    Language-conditioned multi-task policy. Our framework inherits the benefit from diffusion models such that it can condition on language or other modalities to deploy on various tasks with a single set of weights, with negligible performance change compared to per-task policy training.
@@ II Related work
End-to-end Imitation Learning. Significant advancements have been achieved in end-to-end imitation learning [ze20243d, chi2023diffusion, shridhar2023perceiver, goyal2024rvt, du2024learning, fu2024mobile], which involves directly mapping raw sensory observations to action predictions. This field can be broadly divided into two categories: discriminative models and generative models. The discriminative model aims to directly regress the action output given the observations from the collected demonstration data [shridhar2023perceiver, goyal2024rvt]. Representative work PerAct [shridhar2023perceiver] encodes the RGBD image of the entire scene as a 3D volume. With additional language conditioning, it achieves promising results with a low number of demonstrations. RVT [goyal2023rvt, goyal2024rvt] further enhances its efficiency by projecting the 3D scene cloud into orthogonal virtual views. However, the discriminative model often struggles with challenging action prediction tasks with multimodal distributions, sequential actions, or high precision. Recent works [chi2023diffusion, ze20243d, ke20243d, yan2024dnact] leverage diffusion model to learn an action distribution and synthesize action conditioning on the raw sensory input. Despite promising results, this work is limited by the need for extensive demonstration data, limited cross-embodiment generalization, and difficulty adapting to new scenarios.
@@
Imitation Learning from Video. Recent advancements have shifted from relying on explicit robot action data to learning from visual demonstrations, which can be performed by either humans or robots [bahl2022human, du2024learning, pari2021surprising, nair2022r3m, sivakumar2022robotic, kurutach2018learning, karnan2022adversarial, baker2022video, ryu2022equivariant, sharma2019third, wen2022you]. Some recent approaches [vecerik2024robotap, sieb2020graph] extract object state information from video demonstrations and frame the manipulation problem as action replay via low-level visual servoing primitive. Although these methods show promise, reasoning solely over pick and place ignores intermediate actions, preventing applications to tasks where intermediate action needs to satisfy certain constraints. Other methods [wen2023any, xu2024flow, lin2024flowretrieval] use optical flow or point tracking as intermediate representations from video demonstrations, with additional policy networks to output actions. However, these policies rely on robot action demonstrations tied to the same embodiment, restricting cross-embodiment transfer. Another line of work synthesizes videos of desired task executions and determines action by learning an inverse-dynamics model [du2024learning] or extracting target object poses using optical flow [ko2023learning]. Instead of synthesizing videos, which can be computationally intensive and noisy, we directly generate the object pose trajectory end-to-end.
@@
Object-Centric Manipulation. A line of research leverages object-centric representations to reason about visual scenes in a modular way to promote policies prioritizing task-relevant factors while reducing the influence of irrelevant visual distractions [zhu2023learning, wen2022you]. Varying forms of object centric representations have been explored, including object pose [wen2022you, migimatsu2020object, morgan2021vision, morgan2022complex, vitiello2023one], object detection/segmentation [valassakis2022demonstrate, devin2018deep, wang2019deep, zhu2023viola], point tracking [xu2024flow, yuan2024general, huang2024rekep, zhu2024vision, zhang2023flowbotpp, seita2023toolflownet, vecerik2024robotap, gao2021kpam, sundaresan2023kite], neural implicit field [simeonov2022neural, urain2023se, weng2023neural, jauhri2023learning], or object-centric latent features extracted from visual foundation models [di2024dinobot, di2024keypoint]. Among these, it is common to combine object-centric perception with conventional planning [morgan2021vision], or only focus on learning complex policies for the last-inch manipulation [wen2022you, valassakis2022demonstrate, di2024dinobot]. Therefore, they typically rely on manually programmed planning constraints for certain tasks where intermediate sequential action becomes critical [gao2021kpam, manuelli2019kpam, huang2024rekep], and non-trivial to scale to multiple tasks under a single set of policy/hyper-parameters. Moreover, the typical treatment of a constant rigid transform between gripper and object results in open-loop execution and thus lacks reaction to dynamic uncertainties [sundaresan2023kite, gao2021kpam, manuelli2019kpam]. Our framework inherits the prominent generalization properties of object-centric representations by building upon the recent visual foundation models, including object pose estimation and segmentation. We additionally automate the task learning for the complete episode to minimize human intervention by reasoning over sequential object pose trajectories. Closed-loop feedback during robot execution further enhances robustness to dynamic uncertainties when dealing with high-precision tasks.
@@
Figure 2: Overview. During training, we extract object pose trajectories from demonstration RGBD videos (e.g., collected with an iPhone), which are independent of the embodiment. Using these extracted trajectories, we train a diffusion model to generate future object trajectories and determine task completion based on current and past poses. During task execution, the task-relevant object is constantly tracked, and its pose is forwarded to the trajectory diffusion model to predict the object’s future trajectory in SE(3), which leads to task accomplishment. Finally, we convert the generated trajectories into embodiment-agnostic action plans for closed-loop manipulation.
@@ III Problem Setup
We aim to learn a policy $\pi$ that maps the observation $O_{t}$ to the robot’s action $A_{t:t+h}$, where $h$ is the prediction horizon. A handful of demonstrations are provided for learning, which can be performed by either a robot or a human and recorded as RGBD videos. The policy focuses on prehensile manipulation tasks involving rigid objects. The observation $O_{t}$ represents the robot’s sensory data, including RGBD images. For the action space, we utilize positional control, meaning $A_{t}\in SE(3)$ is specified by end-effector poses. In the case of single-policy for multi-tasks, task descriptions are provided to specify task-relevant objects and desired goals to distinguish each task.
@@ IV Approach
Fig. 2 illustrates our imitation learning approach that utilizes object pose trajectories as an intermediate representation to learn the manipulation policy. First, we extract object pose trajectories from demonstration videos, which can be performed by either a robot or a human. Next, with the extracted object pose trajectories, we train a diffusion model to generate future object trajectories conditioning on the current and past object poses. In the case of a single-policy multi-task setting, the diffusion model can additionally condition on a language embedding extracted from the task description to distinguish each task. Finally, we explain how to transform the generated object trajectories into embodiment-agnostic action plans for closed-loop manipulation. Unlike end-to-end imitation learning methods that tie the grasping pose with subsequent motion, our object-centric approach is independent of specific robot hardware or grasp pose. This flexibility enables our method to integrate seamlessly with any online or offline grasp selection technique.
@@ IV-A Demonstration Video Parsing
We opt for object pose trajectories as the object-centric representation, which depicts the full SE(3) state information of rigid objects throughout the episode. Unlike previous methods that derive SE(3) poses from dense point flow [vecerik2024robotap], which can introduce redundancy and noise, directly reasoning about object poses enhances accuracy, as demonstrated in our experiments. Additionally, the compact nature of object poses improves compatibility with diffusion model training and boosts learning efficiency. To learn a policy from the demonstration video, the key step is to extract task-relevant objects and their pose trajectories.
Without loss of generality, we consider a single-arm manipulation scenario where only one object is manipulated at any given time. As a result, the task involves a set of objects $V=\{v_{a},v_{b}\}$, consisting of a graspable source object $v_{a}$ and a target object $v_{b}$. For example, in a water-pouring task, the graspable object might be a kettle, and the target object might be a mug. Given the demonstration video with length $l$, the goal is to extract pose trajectory $\{\hat{T}_{a}\}_{0}^{l}$ and $\{\hat{T}_{b}\}_{0}^{l}$ where each corresponding to object $v_{a}$ and $v_{b}$. These trajectories describe the task-specific object-centric motions in 3D and thus implicitly encode the planning constraints throughout the task, including those intermediates, such as keeping the kettle upright to prevent spilling until it approaches the mug. In practice, we apply a zero-shot 6D object pose estimation method to extract the pose trajectory from the video (Sec. IV-C) for both the source and target objects. We then treat the source and target objects as two nodes in a scene graph and convert all source object poses into the target object’s frame $\hat{\tau}=\{\hat{T}^{b}_{a}\}_{0}^{l}$. This transforms multiple demonstration trajectories into a canonicalized space, regardless of their absolute configurations or backgrounds, allowing the demonstration data to be collected in various environments.
Training directly on the dense, continuous pose trajectory can be inefficient and unstable due to the noisy, shaky motions from human hands or random pauses in the video. To address this, we employ a simple heuristic strategy to subsample keyframes. Specifically, a frame is added if the relative velocity is zero (indicating a change in direction) or if it exceeds a certain distance threshold from the previous keyframe, measured in both translation and rotation. More details on keyframe selection can be found in [shridhar2023perceiver].

### v2

@@ Abstract
We introduce SPOT, an object-centric imitation learning framework. The key idea is to capture each task by an object-centric representation, specifically the SE(3) object pose trajectory relative to the target. This approach decouples embodiment actions from sensory inputs, facilitating learning from various demonstration types, including both action-based and action-less human hand demonstrations, as well as cross-embodiment generalization. Additionally, object pose trajectories inherently capture planning constraints from demonstrations without the need for manually-crafted rules. To guide the robot in executing the task, the object trajectory is used to condition a diffusion policy. We systematically evaluate our method on simulation and real-world tasks. In real-world evaluation, using only eight demonstrations shot on an iPhone, our approach completed all tasks while fully complying with task constraints. Project page: https://nvlabs.github.io/object_centric_diffusion
@@ I Introduction
Learning from demonstrations is an effective method for acquiring complex robotic manipulation skills. This approach simplifies learning vision-based manipulation policies by training directly on demonstration data, thus avoiding the need for intricate hand-crafted rules or laborious robot self-exploration in real-world settings. However, developing a structural representation that captures task-relevant information and remains resilient to environmental perturbations continues to be a significant challenge.
Previous work [reed2022generalist, jang2022bc, ahn2022can] has utilized raw RGB inputs for end-to-end visuomotor policy training. However, the absence of explicit 3D representation hinders generalization to various viewing angles and out-of-distribution spatial configurations. Instead, RVT [goyal2023rvt] and PerAct [shridhar2023perceiver] use the entire 3D scene as input represented by point clouds or 3D volumes. While these methods consider the overall 3D scene structure, they include excessive redundant information, necessitating extensive demonstration data collection. Additionally, the tight coupling of observation and action pairs complicates deployment across different camera setups, or learning from action-less data like human hand demonstrations.
Object-centric methods address these problems by extracting object-centric information from visual observation, such as object detection [zhu2023viola], 6D object poses [wen2022you], and 2D or 3D flows [zhu2024vision, xu2024flow, yuan2024general], to serve as inputs for downstream policies. By leveraging visual foundation models, these methods exhibit impressive generalization capabilities in new environments and demonstrate strong data efficiency. Nevertheless, most works do not model the complete object pose trajectories or focus only on the last-inch manipulation policy. This oversight necessitates extra human effort to manually create rules for constrained planning, especially in tasks where intermediate actions are crucial for success. For example, in tasks like pouring water or serving a plate, the container must remain upright throughout to prevent spillage. Such constraints are non-trivial to automatically learn with existing object-centric frameworks.
@@
To address these issues, we introduce SPOT, an object-centric diffusion policy framework that combines diffusion policies and an object-centric representation for manipulation. As shown in Fig. 1, it uses the SE(3) object-centric trajectory relative to the target as the representation. This approach decouples embodiment action from sensory input, facilitating learning from diverse embodiments and demonstration data, whether action-based or action-less. In particular, in contrast to the standard diffusion policy [chi2023diffusion, ze20243d] which directly produces robot end-effector actions, our diffusion model learns to synthesize target SE(3) object-centric pose trajectories along the horizon, conditioned on the current observed object pose. The training data is obtained by extracting such object-centric information from demonstration videos. For robot execution, the action is computed from the task space to transport the object following the target synthesized trajectory. By recurrently switching between future trajectory synthesis and trajectory following, we achieve closed-loop control with enhanced robustness under dynamic uncertainty, such as in-hand object motion during grasping or placement. Unlike previous methods that focus solely on the last inch of manipulation, our approach trains on the pose trajectory data over the complete horizon, eliminating the need for manually-crafted global planning rules before the last-inch stage. It also leverages data scaling laws when large amounts of demonstration data are available while maintaining data efficiency in few-shot demonstration scenarios.
We evaluate the approach on RLBench [james2020rlbench], where SPOT outperforms all baselines in high-precision and long-horizon tasks in a challenging single-camera setting. In real-world tasks, with just 8 demonstrations shot on an iPhone, the resulting policies complete all tasks while fully complying with task constraints. We will make all datasets and code available so others can reproduce and expand upon this work. In summary, our framework has the following desired properties:
@@
    Flexibility on diverse demonstration data. Our framework can learn from action-based robotic demonstrations and action-less data from human demonstrations. This flexibility enables easier data collection using readily-accessible setups, such as a handheld iPhone, without the need for robotic arms, as our experiments demonstrate.
@@
    Autonomy on complex task constraints. Our framework considers the entire sequential object pose trajectory. This approach allows it to automatically learn planning constraints from demonstrations, eliminating the need for manually-crafted rules.
@@
    Generalization. Our object-centric representation allows us to ignore task-irrelevant information from the background and thus generalizes robustly to varying lighting conditions, scene setup, and sensory configurations.
@@
    Language-conditioned multi-task policy. Our framework inherits the benefits of diffusion models such that it can condition on language or other modalities to deploy on various tasks with a single set of weights, with negligible performance change compared to per-task policy training.
@@ II Related work
End-to-end Imitation Learning. Significant advancements have been achieved in end-to-end imitation learning [ze20243d, chi2023diffusion, shridhar2023perceiver, goyal2024rvt, du2024learning, fu2024mobile], which involves directly mapping raw sensory observations to action predictions. This field can be broadly divided into two categories: discriminative models and generative models. The discriminative model aims to directly regress the action output given the observations from the collected demonstration data [shridhar2023perceiver, goyal2024rvt]. Representative work PerAct [shridhar2023perceiver] encodes the RGBD image of the entire scene as a 3D volume. With additional language conditioning, it achieves promising results with a low number of demonstrations. RVT [goyal2023rvt, goyal2024rvt] further enhances its efficiency by projecting the 3D scene cloud into orthogonal virtual views. However, the discriminative model often struggles with challenging action prediction-tasks with multimodal distributions, sequential actions, or high precision. Recent works [chi2023diffusion, ze20243d, ke20243d, yan2024dnact] leverage diffusion models to learn an action distribution and synthesize action conditioned on the raw sensory input. Despite promising results, this work is limited by the need for extensive demonstration data, limited cross-embodiment generalization, and difficulty adapting to new scenarios.
@@
Imitation Learning from Video. Recent advancements have shifted from relying on explicit robot action data to learning from visual demonstrations, which can be performed by either humans or robots [bahl2022human, du2024learning, pari2021surprising, nair2022r3m, sivakumar2022robotic, kurutach2018learning, karnan2022adversarial, baker2022video, ryu2022equivariant, sharma2019third, wen2022you]. Some recent approaches [vecerik2024robotap, sieb2020graph] extract object state information from video demonstrations and frame the manipulation problem as action replay via low-level visual servoing primitives. Although these methods show promise, reasoning solely on action primitives like picking and placing ignores intermediate actions, preventing applications to tasks where intermediate action needs to satisfy certain constraints. Other methods [wen2023any, xu2024flow, lin2024flowretrieval] use optical flow or point tracking as intermediate representations from video demonstrations, with additional policy networks to output actions. However, these policies rely on robot action demonstrations tied to the same embodiment, restricting cross-embodiment transfer. Another line of work synthesizes videos of desired task executions and determines actions by learning an inverse-dynamics model [du2024learning] or extracting target object poses using optical flow [ko2023learning]. Instead of synthesizing videos, which can be computationally intensive and noisy, we directly generate the object pose trajectory end-to-end.
@@
Object-Centric Manipulation. A line of research leverages object-centric representations to reason about visual scenes in a modular way to promote policies prioritizing task-relevant factors while reducing the influence of irrelevant visual distractions [zhu2023learning, wen2022you]. Varying forms of object centric representations have been explored, including object pose [wen2022you, migimatsu2020object, morgan2021vision, morgan2022complex, vitiello2023one], object detection/segmentation [valassakis2022demonstrate, devin2018deep, wang2019deep, zhu2023viola], point tracking [xu2024flow, yuan2024general, huang2024rekep, zhu2024vision, zhang2023flowbotpp, seita2023toolflownet, vecerik2024robotap, gao2021kpam, sundaresan2023kite], neural implicit fields [simeonov2022neural, urain2023se, weng2023neural, jauhri2023learning], or object-centric latent features extracted from visual foundation models [di2024dinobot, di2024keypoint]. Among these efforts, it is common to combine object-centric perception with conventional planning [morgan2021vision], or only focus on learning complex policies for last-inch manipulation [wen2022you, valassakis2022demonstrate, di2024dinobot]. These approaches often depend on manually-programmed planning constraints for tasks where intermediate sequential actions are critical [gao2021kpam, manuelli2019kpam, huang2024rekep], making it difficult to scale to multiple tasks with a single policy and set of hyperparameters. Moreover, the typical treatment of a constant rigid transform between gripper and object results in open-loop execution and lacks reactivity to dynamic uncertainties [sundaresan2023kite, gao2021kpam, manuelli2019kpam]. Our framework inherits the prominent generalization properties of object-centric representations by building upon recent visual foundation models, including object pose estimation and segmentation. We additionally automate task learning for the complete episode and minimize human intervention by reasoning over sequential object pose trajectories. Closed-loop feedback during robot execution further enhances robustness to dynamic uncertainties when dealing with high-precision tasks.
@@
Figure 2: Overview. During training, we extract object pose trajectories from demonstration RGBD videos (e.g., collected with an iPhone), which are independent of the embodiment. Using these extracted trajectories, we train a diffusion model to generate future object trajectories and determine task completion based on current and past poses. During task execution, the task-relevant object is constantly tracked, and its pose is forwarded to the trajectory diffusion model to predict the object’s future trajectory in SE(3) that leads to task accomplishment. Finally, we convert the generated trajectories into embodiment-agnostic action plans for closed-loop manipulation.
@@ III Problem Setup
We aim to learn a policy $\pi$ that maps the observation $O_{t}$ to the robot’s action $A_{t:t+h}$, where $h$ is the prediction horizon. A handful of demonstrations are provided for learning, which can be performed by either a robot or a human and recorded as RGBD videos. The policy focuses on prehensile manipulation tasks involving rigid objects. The observation $O_{t}$ represents the robot’s sensory data, including RGBD images. For the action space, we utilize positional control, meaning $A_{t}\in SE(3)$ is specified by end-effector poses. In the case of a single multi-task policy, task descriptions are provided to specify task-relevant objects and desired goals.
@@ IV Approach
Fig. 2 shows our approach, which uses object pose trajectories as an intermediate representation for manipulation policy learning. We extract object trajectories from demonstration videos that can be performed by either a robot or a human. With the extracted object pose trajectories, we train a diffusion model to predict future object trajectories conditioned on the object pose history. For multi-task settings, the model can condition on a task description embedding. The generated trajectories are then converted into embodiment-agnostic action plans for closed-loop manipulation. Unlike end-to-end methods that tie grasp poses to subsequent motion, our object-centric approach is hardware-independent, allowing seamless integration with any grasp selection technique.
@@ IV-A Demonstration Video Parsing
We opt for object pose trajectories as the object-centric representation, which depicts the full SE(3) state information of rigid objects throughout the episode. Previous methods that derive SE(3) poses from dense point flow [vecerik2024robotap] can introduce redundancy and noise; instead, directly reasoning about object poses enhances accuracy, as demonstrated in our experiments. Additionally, the compact nature of object poses improves compatibility with diffusion model training and boosts learning efficiency. To learn a policy from the demonstration video, the key step is to extract task-relevant objects and their pose trajectories.
Without loss of generality, we consider a single-arm manipulation scenario where only one object is manipulated at any given time. As a result, the task involves a set of objects $V=\{v_{a},v_{b}\}$, consisting of a graspable source object $v_{a}$ and a target object $v_{b}$. For example, in a water-pouring task, the graspable object might be a kettle, and the target object might be a mug. Given the demonstration video with length $l$, the goal is to extract pose trajectory $\{\hat{T}_{a}\}_{0}^{l}$ and $\{\hat{T}_{b}\}_{0}^{l}$ corresponding to object $v_{a}$ and $v_{b}$, respectively. These trajectories describe the task-specific object-centric motions in 3D and thus implicitly encode the planning constraints throughout the task, including intermediates constraints such as keeping the kettle upright to prevent spilling until it approaches the mug. In practice, we apply a zero-shot 6D object pose estimation method [wen2024foundationpose] to extract the pose trajectory from the video (Sec. IV-C) for both the source and target objects. We then treat the source and target objects as two nodes in a scene graph and convert all source object poses into the target object’s frame $\hat{\tau}=\{\hat{T}^{a}_{b}\}_{0}^{l}$. This transforms multiple demonstration trajectories into a canonical space regardless of their absolute configurations or backgrounds, allowing demonstration data to be collected in various environments.
Training directly on the dense, continuous pose trajectory can be inefficient and unstable due to noisy, shaky motions from human hands in the video. To address this, we adopt the keyframe selection approach described in PerAct [shridhar2023perceiver], which is summarized as follows: a frame is added if the relative velocity is zero (indicating a change in direction) or if it exceeds a certain distance threshold from the previous keyframe, measured in both translation and rotation.